# How Hungry is AI? Benchmarking Energy, Water, and Carbon Footprint of LLM Inference

### v4

@@ 5.2 Water and Carbon Emissions
Figure 3 showcases the water consumption and carbon emissions of models across all prompt sizes. The most resource-efficient systems, including GPT-4.1 nano, LLaMA-3.2 1B, and LLaMA-3.2 3B, maintain carbon emissions below 0.3 grams per query while using less than 2 milliliters of water across all input lengths.
By contrast, models such as DeepSeek-R1, DeepSeek-V3, o3, and GPT-4.5 exhibit substantially larger environmental footprints across all input sizes. DeepSeek-R1 consistently emits over 14 grams of carbon dioxide and consumes more than 150 milliliters of water per query. For reference, this is equivalent to driving 50 meters in a gasoline-powered car and using two-thirds of a standard water cup. These figures suggest that environmental impacts are shaped not only by model architecture but also by deployment strategies and regional infrastructure conditions. In particular, the elevated emissions and water usage observed in DeepSeek models likely reflect inefficiencies in their data centers, including higher PUE and suboptimal cooling technologies.
@@
6 GPT-4o Case Study
@@ 6.3 Estimated 2025 Annual Water Footprint of GPT-4o Inference
As showcased in Figure 4, we translate estimated cooling and infrastructure-related water usage into real-world benchmarks. Based on scaled inference volumes, GPT-4o’s annual water consumption is projected to be between 1,334,991 kiloliters (kL) and 1,579,680 kL. These quantities are roughly equivalent to filling over 500 Olympic-sized swimming pools (1,250,000 kL). Importantly, this consumption refers to evaporated freshwater permanently removed from local ecosystems rather than recycled. GPT-4o alone is responsible for evaporating an amount of freshwater equivalent to the annual drinking needs of almost 1.2 million people.
@@ 7.1 The Critical Role of Infrastructure in AI Sustainability
Our findings indicate that infrastructure is a crucial determinant of AI inference sustainability. While model design enhances theoretical efficiency, real-world outcomes can substantially diverge based on deployment conditions and factors such as renewable energy usage and hardware efficiency. For instance, GPT-4o mini, despite its smaller architecture, consumes approximately 20% more energy than GPT-4o on long queries due to reliance on older A100 GPU nodes. Similarly, DeepSeek models exhibit disproportionately high water footprints, not solely due to model characteristics but due to data center inefficiencies. These observations suggest that true sustainability will depend on integrating more efficient hardware, sustainable cooling strategies, renewable energy sourcing, evaluation practices, and deployment infrastructures.
@@ 8 Conclusion, Limitations, and Future Work
This paper introduces the first large-scale, infrastructure-aware framework for benchmarking the environmental footprint of LLM inference, integrating API performance, environmental multipliers, and statistical inference to assess energy, water, and carbon costs under real-world conditions. By applying cross-efficiency DEA, we contextualize environmental impact in terms of functional performance, revealing that eco-efficiency hinges not only on model design but also on infrastructure. Our GPT-4o case study emphasizes the Jevons Paradox: As AI becomes cheaper and faster, total usage expands, intensifying environmental strain despite gains in per-query efficiency. Without structural shifts in how LLMs are designed and deployed, these invisible costs will continue to rise, threatening to offset the societal benefits that made these systems valuable in the first place. This work establishes a standardized, scalable framework for benchmarking the environmental footprint of LLM inference in real-world data center deployments, providing a basis for transparent, infrastructure-aware sustainability assessment and future regulation.
Our work inherits certain limitations that we acknowledge; we avoid overstating model-specific footprints by conservatively including only the energy drawn by actively assigned GPUs. This is due to the lack of means to determine whether unused GPUs’ capacity is reassigned, load-balanced, or left inactive. Isolating non-GPU power consumption was also difficult. We applied a fixed utilization estimate from prior studies, acknowledging that their variation across inference workloads is typically significantly lower than that of GPUs. Where facility-specific data were unavailable, we used regional or national averages for PUE, WUE, and CIF. Moreover, for proprietary models without disclosed size, we classified their scale based on observed API performance. Future work should address these limitations as more detailed telemetry and facility-level reporting become available. Additionally, future studies should also extend beyond text generation to evaluate image, video, and audio generation, which are likely to impose greater environmental costs due to higher computational intensity.
@@ Appendix C Cross-effficiency DEA Results
Before presenting the eco-efficiency results, it is worth noting that GPT-4, GPT-4 Turbo, and LLaMA-3.2 Vision 11B were excluded due to the lack of benchmark results on certain tests due to model limitations. Since cross-efficiency requires complete inputs and outputs, these models could not be fairly evaluated.
As shown in Figure 6, Anthropic’s newest model dominates the eco-efficiency frontier. Claude-3.7 Sonnet scored highest (0.886), combining strong reasoning with an efficient infrastructure footprint. OpenAI’s o4-mini (high) (0.867) and o3-mini (0.840) also performed well, offering solid multi-step reasoning at lower resource cost. These results suggest that downsizing reasoning models can yield substantial sustainability gains with minimal performance trade-offs.
At the opposite end, DeepSeek-R1 (0.058) and DeepSeek-V3 (0.060) had the lowest scores. Despite high intelligence ratings, their energy, water, and carbon demands are disproportionately high, indicating severe infrastructural inefficiencies. Among OpenAI models, GPT-4.1 mini (0.802) and GPT-4o (0.762) were strong performers, balancing intelligence and environmental impact. In contrast, GPT-4.5 ranked among the least efficient, showing that newer architectures do not always yield more sustainable outcomes. LLaMA models clustered near 0.5, limited by weaker reasoning scores. Although resource-efficient, their modest performance kept overall eco-efficiency low.
In summary, eco-efficiency relies on both output quality and environmental cost. Anthropic’s newest model and OpenAI’s smaller reasoning models excel in both areas, while DeepSeek and LLaMa models demonstrate the limitations of concentrating on capability or sustainability alone.

### v5

@@ 5.2 Water and Carbon Emissions
Figure 4 showcases the water consumption and carbon emissions of models across all prompt sizes. The most resource-efficient systems, such as LLaMA-3.2 1B, LLaMA-3.2 3B, LLaMA-3.1-8B, LLaMA-3-8B, and GPT-4.1 nano, emit less than 0.3 gCO2e and consume under 4 mL of water even for long-form prompts, demonstrating exceptional sustainability across scales.
In contrast, large-scale and reasoning models such as o3, DeepSeek-R1 (DS), and DeepSeek-V3 (DS) exhibit substantially higher footprints. DeepSeek-R1 (DS) consumes over 200 mL of water and emits approximately 17 gCO2e per long query, while the same model on Azure consumes only 34 mL and emits 2.5 gCO2e, a reduction of nearly 85%. These figures suggest that environmental impacts are shaped not only by model architecture but also by deployment strategies and regional infrastructure conditions. In particular, the elevated emissions and water usage observed in DeepSeek models likely reflect inefficiencies in their data centers, including higher PUE, suboptimal cooling technologies, and less efficient hardware.
@@
5.3 Validation Against Public Disclosures
Public disclosures of inference-level energy and carbon data remain limited, but a few recent statements provide useful reference points for cross-validation. In June 2025, OpenAI CEO Sam Altman reported that the default ChatGPT model consumed approximately 0.34 Wh per query [68]. Knowing that GPT-4o was the default deployment at that time, this estimate likely corresponds to GPT-4o-level inference. Our framework estimates 0.42 Wh ($\pm$0.13 Wh) for a short GPT-4o prompt (0.37 Wh without datacenter overhead), within 19% of Altman’s figure. Similarly, the results for Mistral Large 2 align closely with Mistral’s published life-cycle assessment (LCA) report [69], which cites approximately 1.14 gCO2e per 400-token query. Our corresponding estimate for 300 tokens (0.82 gCO2e, $\pm$0.10 gCO2e) scales to roughly 1.09 gCO2e when normalized to 400 tokens, showcasing alignment within one standard deviation. Together, these alignments between independent disclosures and our modeled results suggest that the framework reproduces realistic operational conditions for modern LLM inference.
6 GPT-4o Environmental Impact Case Study
@@ 6.3 Estimated 2025 Annual Water Footprint of GPT-4o Inference
As showcased in Figure 5, we translate estimated cooling and infrastructure-related water usage into real-world benchmarks. Based on scaled inference volumes, GPT-4o’s annual water consumption is projected to be between 1,334,991 kiloliters (kL) and 1,579,680 kL. These quantities are roughly equivalent to filling over 500 Olympic-sized pools or to supporting the annual drinking needs of 1.2 million people. Importantly, this consumption refers to evaporated freshwater permanently removed from local ecosystems rather than recycled. GPT-4o alone is responsible for evaporating an amount of freshwater equivalent to the annual drinking needs of almost 1.2 million people.
@@ 6.4 Estimated 2025 Annual Carbon Footprint of GPT-4o Inference
7 GPT-5 Adaptive Model Routing Case Study
The launch of GPT-5 [74] introduced adaptive model routing, a mechanism that allows the system to automatically determine whether to use a fast variant or a more computationally intensive “Thinking” model for complex reasoning tasks. This unification eliminates the need for manual model selection where the model dynamically scales its reasoning effort based on prompt complexity.
However, this adaptability introduces substantial variability in energy consumption across reasoning modes, as shown in Figure 6. For medium-length queries, the average energy consumption ranges from 2.33Wh for minimal reasoning to 17.15Wh for high reasoning, representing a more than seven-fold increase. Despite this variance, GPT-5 remains relatively efficient at lower reasoning levels. For instance, a short, minimal reasoning query consumes only 0.67 Wh, a value comparable to GPT-4o’s 0.42 Wh per short prompt. Conversely, a long, high-reasoning query reaches an average of 33.8 Wh, comparable to the upper bounds observed among the most energy-intensive models analyzed in this study.
These results suggest that while adaptive routing optimizes computational resources by tailoring inference depth to task complexity, it also amplifies the environmental footprint of cognitively demanding prompts. This finding underscores the growing importance of prompt-level efficiency analysis for next-generation LLMs that blend lightweight and high-reasoning architectures within a unified system.
Figure 6: Energy consumption of GPT-5 across query lengths and reasoning modes
@@ 8.1 The Critical Role of Infrastructure in AI Sustainability
Our findings indicate that infrastructure is a crucial determinant of AI inference sustainability. While model design enhances theoretical efficiency, real-world outcomes can substantially diverge based on deployment conditions and factors such as renewable energy usage and hardware efficiency. For instance, GPT-4o mini, despite its smaller architecture, consumes approximately 20% more energy than GPT-4o on long queries due to reliance on older A100 GPU nodes. Similarly, DeepSeek models highlight the profound impact of infrastructure: DeepSeek-R1 and DeepSeek-V3 deployed on DeepSeek’s own servers exhibit water consumption and carbon emissions nearly six times higher than their Azure-hosted counterparts. The Azure deployments benefit from better hardware, more efficient cooling systems, lower carbon intensity, and tighter PUE control, demonstrating that sustainability gains can stem as much from datacenter design as from model optimization. These observations underscore that true AI sustainability will hinge on coordinated progress in hardware efficiency, renewable energy sources, and infrastructure-aware deployment strategies.
@@ 9 Conclusion, Limitations, and Future Work
This paper introduces the first large-scale, infrastructure-aware framework for benchmarking the environmental footprint of LLM inference, integrating API performance, environmental multipliers, and statistical inference to assess energy, water, and carbon costs under real-world conditions. By applying cross-efficiency DEA, we contextualize environmental impact in terms of functional performance, revealing that eco-efficiency hinges not only on model design but also on infrastructure. Our GPT-4o case study emphasizes the Jevons Paradox: As AI becomes cheaper and faster, total usage expands, intensifying environmental strain despite gains in per-query efficiency. Additionally, our GPT-5 case study sheds lights on the importance of prompt-level efficiency and adaptive routing. Without structural shifts in how LLMs are designed, deployed, and used, these invisible costs will continue to rise, threatening to offset the societal benefits that made these systems valuable in the first place. This work establishes a standardized, scalable framework for benchmarking the environmental footprint of LLM inference in real-world data center deployments, providing a basis for transparent, infrastructure-aware sustainability assessment and future regulation.
Our work inherits certain limitations that we acknowledge: we avoid overstating model-specific footprints by conservatively including only the energy drawn by actively assigned GPUs. This is due to the lack of means to determine whether unused GPUs’ capacity is reassigned, load-balanced, or left inactive. Isolating non-GPU power consumption was also difficult. We applied a fixed utilization estimate from prior studies, acknowledging that their variation across inference workloads is typically significantly lower than that of GPUs. Moreover, for proprietary models without disclosed size, we classified their scale based on observed API performance. Future work should address these limitations as more detailed telemetry and facility-level reporting become available. Additionally, future studies should also extend beyond text generation to evaluate image, video, and audio generation, which are likely to impose greater environmental costs due to higher computational intensity.
@@ Appendix C Cross-effficiency DEA Results
Before presenting the eco-efficiency results, it is worth noting that Claude 3.5 Sonnet, Claude 3.5 Haiku, GPT-4, and GPT-4 Turbo were excluded due to the lack of benchmark results on certain tests. Since cross-efficiency requires complete inputs and outputs, these models could not be fairly evaluated.
As shown in Figure 8, OpenAI’s reasoning models dominate the eco-efficiency frontier. o3-mini achieved the highest cross-efficiency score (0.884), closely followed by o1-mini (0.836) and Anthropic’s Claude 3.7 Sonnet (0.825), which combines strong reasoning ability with a relatively modest environmental footprint. GPT-4o (Mar) (0.789) and o3 (0.758) also performed well. These results suggest that downsizing reasoning models can yield meaningful sustainability gains without compromising performance.
At the opposite end, DeepSeek-R1 (0.067) and DeepSeek-V3 (0.059) recorded the lowest efficiency scores. Despite their advanced reasoning capabilities, their high energy, water, and carbon costs indicate significant infrastructural inefficiencies. Their Azure-hosted variants performed better, DeepSeek-R1 (0.539) and DeepSeek-V3 (0.523), yet remained below most OpenAI and Anthropic systems. Among OpenAI models, GPT-4.1 mini (0.580) and GPT-4.1 nano (0.508) balanced output quality and sustainability particularly well. LLaMA models clustered between 0.4 and 0.6, reflecting efficient power use but limited reasoning performance.
In summary, eco-efficiency relies on both output quality and environmental cost. OpenAI’s smaller reasoning models and Claude 3.7 Sonnet strike that balance most effectively, while DeepSeek and LLaMA demonstrate the limitations of concentrating on capability or sustainability alone.